# GD doesn’t make the cut: Three ways that non-differentiability affects neural network training

### v3

@@ Abstract
This paper investigates the distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) designed for differentiable functions. First, we demonstrate significant differences in the convergence properties of NGDMs compared to GDs, challenging the applicability of the extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks. Next, we demonstrate the paradoxical nature of NGDM solutions for $L_{1}$-regularized problems, showing that increasing the regularization penalty leads to an increase in the $L_{1}$ norm of optimal solutions in NGDMs. Consequently, we show that widely adopted $L_{1}$ penalization-based techniques for network pruning do not yield expected results. Additionally, we dispel the common belief that optimization algorithms like Adam and RMSProp perform similarly in non-differentiable contexts. Finally, we explore the Edge of Stability phenomenon, indicating its inapplicability even to Lipschitz continuous convex differentiable functions, leaving its relevance to non-convex non-differentiable neural networks inconclusive. Our analysis exposes misguided interpretations of NGDMs in widely referenced papers and texts due to an overreliance on strong smoothness assumptions, emphasizing the necessity for a nuanced understanding of foundational assumptions in the analysis of these systems.
Machine Learning, ICML
@@
Gradient Descent (GD) and its variants (Duchi et al., 2011; Kingma & Ba, 2014; Lydia & Francis, 2019; McMahan & Streeter, 2010; Shi & Li, 2021; Tieleman et al., 2012; Zhang, 2018) have played a pivotal role in advancing image and language processing over the last two decades. These algorithms have found widespread use in neural network training because the are easy to implement, highly scalable, and iteratively approach stationary points, where the loss function’s gradient is zero (Bertsekas, 1997).
@@
Despite being initially designed for differentiable loss functions, GDs are widely used to minimize the loss functions of non-differentiable neural networks. The loss functions of these non-differentiable neural networks takes a special form – they are differentiable almost everywhere. Consequently, when gradient descents are applied to these non-differentiable neural networks (which we refer to as non-differentiable gradient methods or NGDMs), they seldom encounter non-differentiable points during the training process. This trait, combined with recent findings that NGDMs converge to stationary points under mild regularity conditions (Bolte & Pauwels, 2021; Davis et al., 2020), has led standard references in the field to assume that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of (Stevens et al., 2020), for instance). To emphasize this point, we note that three of the most widely cited surveys on optimization methods for neural networks (Bottou et al., 2018; Le et al., 2011; Ruder, 2016) do not mention non-differentiability, even though the ReLU, arguably the most common activation function in neural networks, is non-differentiable.
@@
Moreover, this assumption has sparked renewed interest in the dynamics of gradient descents amid complex loss landscapes ((Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019) to state a few), with fewer studies explicitly addressing non-differentiability (Bertoin et al., 2021; Bolte & Pauwels, 2021; Davis et al., 2020). Additionally, it has led theoretical papers to develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in (Ahn et al., 2022b), Section 3.1 in (Ma et al., 2022), Section 6 in (Zhang et al., 2022) to state a few).
@@
In our paper, we challenge this assumption, demonstrating that it leads to inaccuracies in several highly referenced papers and texts regarding training dynamics in neural networks. We analyze three different aspects of the training process and make the following contributions:
  • Convergence Analysis of ReLU Networks: We show that the rate of convergence of NGDMs is slower than that of GDs. Our findings strongly suggest that the many claims about rate of convergence made under the assumption of $L-smoothness$ have little bearing on the properties of NGDMs. Furthermore, we dispel the widespread belief that different NDGM variants like RMSProp and Adam largely differ only in their rates of convergence, but converge to roughly the same eventual point in the limit.
@@
  • Solution to the LASSO problem: We show that NGDMs, including variants like RMSProp and NDGM with momentum, do not produce sparse solutions for the LASSO problem, even for the simplest case of the $L_{1}$ penalized linear model, and exhibit different behaviors, contrary to prevalent wisdom. Moreover, NGDMs can yield counterintuitive results, such as producing a solution with a larger $L_{1}$ norm with a larger LASSO penalty.
@@
  • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks. Furthermore, we provide a counter-example to the claim made in (Ma et al., 2022) that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
@@
In a broader context, our analysis emphasizes the critical role the assumptions we make play on our understanding of these complex systems. For NGDMs in particular, our results suggest a need to treat GDs and NGDMs as distinct systems with different dynamics, challenging the common assumption of their similarity.
@@ 1 Preliminaries
In the three sections that follow, we show how disregarding the non-differentiability of the loss function severely limits our understanding of different aspects of training dynamics. Each of these sections is largely self-contained, complete with a literature review, problem setup, and analysis. Throughout this paper, experiments involving deep networks consistently employ convolutional neural network (CNN) architectures akin to VGG16 (Simonyan & Zisserman, 2014). This choice is motivated by the architecture’s widespread usage in the papers referenced throughout our discussions. Our experiments have relatively low computational costs, with the most expensive run completing within 2 hours on a single GPU. The present section addresses background and terminology that is pertinent to all the subsequent sections.
NDGM v/s GD:
The crux of our analyses lies in the structural differences between GDs and NDGMs. For a continuously differentiable loss function, $f(\beta)$, the well-studied GD sequence (Bertsekas, 1997; Boyd et al., 2004) is described by the recursion
@@
where ~$\widetilde{\nabla}g(\tilde{\beta}_{t})=\nabla g(\tilde{\beta}_{t})$ if the gradient exists at $\tilde{\beta}_{t}$, else it is a heuristic measure  [^1] https://rb.gy/g74av. The deep learning literature widely assumes that the dynamics described by the recursions in (1) and (2) are nearly the same.
@@
In this paper, we show the problems with this assumption using a variety of convex loss functions. The majority of our analysis is performed on the penalized single layer neural network regression problem with a non-positive response vector, which (Kumar, 2023) has shown to be convex with with the all-zeros vector as its unique global minimizer. The loss function for this problem is given by
$\begin{split}L_{1}(\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})&=||{\bf{y}}-\max(0,{\bf{Z}}\beta)||^{2}\\ &=+\lambda_{1}||\beta||_{1}+\lambda_{2}||\beta||_{2}^{2},\end{split}$ (3)
where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix whose $i^{th}$ column is given by $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, $\lambda_{2}\geq 0$ is the ridge penalty, and $\lambda_{1}+\lambda_{2}>0$. Throughout this paper, we denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations as $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
@@
Our analysis heavily relies on the observation that for a convex non-differentiable loss functions, the NDGM recursion described in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012). The theoretical properties of this method are well studied in the optimization literature, with a concise overview of key findings available in the lecture notes of (Boyd et al., 2003) and (Tibshirani, 2015). The subgradient method and GD recursions have different properties, and these differences play a key role in our subsequent discussions. We reference the relevant portions from the lecture notes as they arise.
@@
Our formulations in equations (1) and (2) employ a constant learning rate. This choice bears some explaining, for which we first recall some definitions from the optimization literature (see Chapter 1 of (Bertsekas, 1997) for details). In the GD literature, the training regimen described in equation (1) is called a constant step-size regime. In this regime, the GD is initialized using some randomized approach, and iterated with a constant learning rate until the loss stops decreasing. Although constant step-size GDs approach the stationary point, they rarely converge to it unless the loss function possesses desirable properties like $L-smoothness$ (Proposition 1.2.3 in (Bertsekas, 1997)). To ensure provable convergence to the stationary point in more general settings, studies often adopt the diminishing step-size regime. This regime is characterized by the iteration
@@
where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$. Convergence results in this regime are derived under ideal conditions, and provide guarantees that the recursion will reach the stationary point at least in the theoretical limit. In practice, diminishing step-size GDs are implemented with a stopping criteria which determines when the solution produced by the GD is good enough to end training. The combination of the diminishing step-size regime and stopping criteria commonly used in the neural network literature is what we term the “reducing” step-size regime.
In this regime, training unfolds as follows: commencing with a weight of $\beta_{0}$, GD is run for some $N_{0}$ iterations with a learning rate of $\alpha_{0}$, followed by $N_{1}$ iterations with a reduced learning rate of $\alpha_{1}<\alpha_{0}$, and so on, until the engineer is satisfied with the solution. In the reducing step-size regime, training involves chaining together a finite number (say $T$) of constant step-size GD runs, with the $(i+1)^{th}$ run having an initialization of $\beta_{u}$ and a learning rate of $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$. While the theoretical guarantees of the diminishing step-size regime may not be realized, the limiting dynamics of the sequence in this regime are exactly those of a constant step-size GD, with learning rates and weight initializations chosen corresponding to the values at the beginning of the $T^{th}$ run. Accordingly, we exclusively study the constant step-size problem, assuming appropriate weights and learning rates have been chosen.
@@ 2 NDGM and convergence analysis
Recent theoretical work by (Davis et al., 2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions. Unfortunately, the conditions of the diminishing step size regime are rarely satisfied in practice. To illustrate this point, consider the Assumption A.3 in (Davis et al., 2020), which requires that in the diminishing step size regime, the sum of the learning rates should diverge. As was previously stated, this is a standard assumption even in the thoeretical convergence analysis of GDs. However, the prevelant step-size reduction we observed in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (StepLR in PyTorch [^2] https://rb.gy/pub4s). With this decay scheme, the learning rates form a convergent geometric series, thus violating the assumption.
For the differentiable case, we know that even when the learning rates form a convergent geometric series, the GD sequence will come within some vicinity of the optimal solution. Therefore, the question arises if the same holds true for NDGM sequences. More precisely, can we state that the dynamics of NDGM and GD sequences are identical?
@@
The dynamics of the two sequences are not identical, and we demonstrate this by arriving at a contradiction. Suppose the sequence $\{\beta_{k}\}_{k=0}^{\infty}$ obtained by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})=L_{1}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1},0,0.01)$ has the same properties as a GD sequence. Then, the entire NDGM sequence can be characterized using the ”Capture Theorem” as described below
@@
Let $x_{k+1}=x_{k}-\alpha\nabla f(x_{k})$ be a sequence generated by running GD on a continuously differentiable function, $f(x)$ such that $f(x_{k+1})\leq f(x_{k})$. If $x^{*}$ is an isolated local minimum of $f(x)$ in a unit ball around $x^{*}$, and $||x_{0}-x^{*}||<1$, then $||x_{k}-x^{*}||<1$ for all $k>0$.
@@
Result follows by using $\bar{\epsilon}=1$ in Proposition 1.2.5 of (Bertsekas, 1997). ∎
Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, then we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$. To validate this claim, we first generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$. Next, we generate a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$. Finally, we generate the sequence $\{\beta_{t}\}_{t=0}^{\infty}$, by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with the entries in $\beta_{0}$ uniformly sampled on $[-1,1]$. As seen in Figure 1, although the loss function is a monotone decreasing function of the iteration number, the infinity norm of the NDGM sequence eventually takes values greater than 1, hence establishing the contradiction. The non-capturing described in Figure 1 is to be expected; as stated in Section 1 of (Boyd et al., 2003), ”Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase”.
Figure 1: Plot of the loss function ($L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$), and infinity norm ($||\beta_{k}||_{\infty}$) of the parameter vector as a function of the iteration number ($k$). Although the loss function is monotone decreasing, the infinity norm of the weights does not ”capture”, hence demonstrating that the dynamics of NDGM and GD are not identical
@@
Given that the dynamics of GDs and NDGMs differ, the question arises what assumptions are appropriate for the study of convergence analysis in neural network training. Majority of the analyses in the literature ((Bottou et al., 2018; Zhang et al., 2022; Défossez et al., 2020; You et al., 2019) to state a few) assume that the loss function, $f(\beta)$ is continuously differentiable, and is $L-smooth$, i.e., there exists a constant $L$, such that
@@
for all $x$, $y$ in the domain of $f$. However, this assumption breaks down for non-differentiable functions, where the gradient does not even exist at various points in the domain. This raises the question if the assumption can be salvaged by confining its validity to a local context around the local optimal solution.
Unfortunately, even this cannot be done; as stated in (Laurent & Brecht, 2018), “local minima of ReLU networks are generically nondifferentiable. They cannot be waved away as a technicality, so any study of the loss surface of such network must invoke nonsmooth analysis”. Moreover, assuming that the loss function is $L-smooth$ grossly over-estimates the speed of convergence of NDGMs. To illustrate this, consider that in order to achieve an error level of $\epsilon$, GD applied to an $L-smooth$ convex loss function, $f$, would require $O(1/\epsilon)$ steps for convergence. In contrast, NDGM applied to a convex non-differentiable loss function, $g$, satisfying the Lipschitz condition
@@
would take $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of (Tibshirani, 2015) for the result, and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$). Thus, it is likely that the vast literature on the convergence analysis of GDs for loss functions satisfying (5) have little bearing on the dynamics of convergence in modern non-smooth neural networks. More generally, we expect the speed of convergence for smooth neural networks to be faster than that of non-smooth neural networks.
@@
We conduct two sets of experiments on CNNs of varying depths to validate our intuition in a more general context (see the Appendix A for architecture details). In each set, we perform two identical simulations for every network, with the only difference being the choice of activation function: one simulation uses the the non-smooth ReLU activation function, while the other uses the smooth and asymptotically equivalent GELU activation function (Hendrycks & Gimpel, 2016). The first set of experiments excludes batch normalization, while the second includes it.
Figure 2 shows that without batch normalization, the gap in convergence speeds between smooth and non-smooth networks widens as network depth increases, aligning with our expectations. This trend is further substantiated by (Hendrycks & Gimpel, 2016), who demonstrate accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper). However, with batch normalization, the difference between ReLU and GELU training curves becomes indistinguishable, emphasizing its importance in the training process. We conclude that the choice of activation function is the primary factor contributing to the difference in convergence speeds between ReLU and GELU curves with increasing depth, as no other principled reason appears evident.
Figure 2: Training curves comparing smooth and non-smooth CNNs (from left to right) of depth 11, 13, and 19, respectively. The top row shows validation losses without batch normalization, while the bottom row includes batch normalization.
@@
As stated in (Bottou, 2010), the assumption of $L-smoothness$ ensures that the gradient of the loss function does not change too rapidly with respect to the parameter vector. However, this assumption which does not hold for ReLU networks. To illustrate this, consider the 1-D case with $q(x)=\text{max}(0,x)$. In this case, $\lim_{h\rightarrow 0^{+}}(1/2h)(\nabla q(h)-\nabla q(-h))=\infty$, indicating that the gradient changes arbitrarily quickly at the origin. This observation necessitates the use of non-smooth analysis for studying the loss functions of ReLU networks, as detailed in (Laurent & Brecht, 2018). The key intuition about non-differentiable functions is that they have kinks corresponding to non-differentiable points, and local minima frequently lie at these kinks. As a result, gradient descents which rely on smoothness assumptions cannot access these points of interest. This realization serves as one of the primary motivations for the development and application of second-order cone programs for convex non-smooth problems (Boyd et al., 2004)  [^3] See 26:30 to 29:00 in https://rb.gy/ab0aeh. The inability of gradient descent methods to effectively maneuver non-smooth functions and the associated pathologies form a recurring theme throughout the remaining sections of this paper. In a subsequent section on the Edge of Stability, we show that insights based on the assumption of $L-smoothness$ are not universally applicable even across all Lipschitz continuous, convex, differentiable neural networks. $L-smoothness$ is one of the strongest assumptions after convexity, guaranteeing the existence and boundedness of the first and second derivatives—criteria unmet by most loss functions in neural network literature.

### v4

@@ Abstract
This paper critically examines the fundamental distinctions between gradient methods applied to non-differentiable functions (NGDMs) and classical gradient descents (GDs) for differentiable functions, revealing significant gaps in current deep learning optimization theory. We demonstrate that NGDMs exhibit markedly different convergence properties compared to GDs, strongly challenging the applicability of extensive neural network convergence literature based on $L-smoothness$ to non-smooth neural networks. Our analysis reveals paradoxical behavior of NDGM solutions for $L_{1}$-regularized problems, where increasing regularization counterintuitively leads to larger $L_{1}$ norms of optimal solutions. This finding calls into question widely adopted $L_{1}$ penalization techniques for network pruning. We further challenge the common assumption that optimization algorithms like RMSProp behave similarly in differentiable and non-differentiable contexts. Expanding on the Edge of Stability phenomenon, we demonstrate its occurrence in a broader class of functions, including Lipschitz continuous convex differentiable functions. This finding raises important questions about its relevance and interpretation in non-convex, non-differentiable neural networks, particularly those using ReLU activations. Our work identifies critical misunderstandings of NDGMs in influential literature, stemming from an overreliance on strong smoothness assumptions. These findings necessitate a reevaluation of optimization dynamics in deep learning, emphasizing the crucial need for more nuanced theoretical foundations in analyzing these complex systems.
1 Introduction
@@
Gradient Descent (GD) and its variants (Duchi et al., 2011; Kingma and Ba, 2014; Lydia and Francis, 2019; McMahan and Streeter, 2010; Shi and Li, 2021; Tieleman et al., 2012; Zhang, 2018) have been instrumental in advancing image and language processing over the last two decades. These algorithms have gained widespread adoption in neural network training due to their ease of implementation, high scalability, and ability to iteratively approach stationary points where the loss function’s gradient vanishes (Bertsekas, 1997).
@@
However, an important consideration has emerged in the field: despite being originally designed for differentiable loss functions, GDs are routinely applied to minimize loss functions of non-differentiable neural networks. These non-differentiable loss functions possess a unique characteristic—they are differentiable almost everywhere. Consequently, when gradient descents are applied to these non-differentiable neural networks (which we term non-differentiable gradient methods or NGDMs), they rarely encounter non-differentiable points during training. This property, coupled with recent findings suggesting NGDMs converge to stationary points under mild regularity conditions (Bolte and Pauwels, 2021; Davis et al., 2020), has led to a widespread assumption in the field: that non-differentiability in neural networks is inconsequential (see Section 6.1.5 of Stevens et al. (2020), for instance). Notably, three of the most widely cited surveys on optimization methods for neural networks (Bottou et al., 2018; Le et al., 2011; Ruder, 2016) do not discuss non-differentiability, despite the ReLU—arguably the most ubiquitous activation function in neural networks—being non-differentiable.
@@
This assumption has not only rekindled interest in gradient descent dynamics within complex loss landscapes (Ahn et al., 2022b; Dauphin et al., 2014; Reddi et al., 2019), but has also led to a trend where theoretical papers develop results for continuously differentiable functions and then claim universality by running simulations on non-differentiable neural networks (Experiment 6 in Ahn et al. (2022b), Section 3.1 in Ma et al. (2022), Section 6 in Zhang et al. (2022)).
@@
Our paper examines this fundamental assumption, demonstrating that it leads to significant inaccuracies in several highly referenced papers and texts regarding neural network training dynamics. We analyze three critical aspects of the training process, making the following key contributions:
  • Convergence Analysis of ReLU Networks: We show that NGDMs converge more slowly than GDs, suggesting that many convergence rate claims made under the assumption of $L-smoothness$ have limited applicability to NGDM properties.
@@
  • Solutions to the LASSO problem: We demonstrate that NGDMs do not produce sparse solutions for the LASSO problem, even in the simplest case of the $L_{1}$ penalized linear model. This finding contrasts with prevalent understanding. Moreover, we show that NGDMs can yield unexpected results, such as producing solutions with larger $L_{1}$ norms when using larger LASSO penalties. Importantly, we prove that variants like RMSProp and NDGM with momentum behave differently from vanilla NDGM, challenging the common assumption that these algorithms converge to similar solutions in non-differentiable settings.
@@
  • The Edge of Stability: We demonstrate that the edge of stability conjecture does not hold, even for gradient descents on all convex differentiable neural networks. Furthermore, we provide a counter-example to the claim made in Ma et al. (2022) that subquadratic behavior around the local minimum is responsible for the Edge of Stability phenomenon.
@@
In a broader context, our analysis underscores the importance of the assumptions we make in understanding these complex systems. For NGDMs in particular, our results suggest a need to reconsider how we approach these methods, challenging the common assumption of their similarity to GDs and calling for a more nuanced theoretical foundation in deep learning optimization.
@@ 2 Preliminaries
In the following three sections, we demonstrate how disregarding the non-differentiability of the loss function significantly limits our understanding of various aspects of training dynamics. Each section is largely self-contained, including its own literature review, problem setup, and analysis. Throughout this paper, our experiments with deep networks consistently employ convolutional neural network (CNN) architectures similar to VGG16 (Simonyan and Zisserman, 2014), chosen for its widespread use in the papers we reference. Our experiments have relatively low computational demands, with the most expensive run completing within 2 hours on a single GPU. This section introduces background and terminology relevant to all subsequent sections.
NDGM vs. GD:
The core of our analyses lies in the structural differences between GDs and NDGMs. For a continuously differentiable loss function, $f(\beta)$, the well-studied GD sequence (Bertsekas, 1997; Boyd et al., 2004) is described by the recursion
@@
where ~$\widetilde{\nabla}g(\tilde{\beta}_{t})=\nabla g(\tilde{\beta}_{t})$ if the gradient exists at $\tilde{\beta}_{t}$, else it is a heuristic measure  [^1] https://rb.gy/g74av. The deep learning literature widely assumes that the dynamics described by the recursions in (1) and (2) are nearly identical.
@@
We demonstrate the issues with this assumption using various convex loss functions. Most of our analysis focuses on the penalized single-layer neural network regression problem with a non-positive response vector, which Kumar (2023) has shown to be convex with the all-zeros vector as its unique global minimizer. The loss function for this problem is:
$L_{1}($ $\beta;{\bf{Z}},{\bf{y}},\lambda_{1},\lambda_{2})=$
$\|{\bf{y}}-\max(0,{\bf{Z}}\beta)\|^{2}+\lambda_{1}\|\beta\|_{1}+\lambda_{2}\|\beta\|_{2}^{2},$ (3)
where $\beta$ is a $P\times 1$ parameter vector, ${\bf{Z}}$ is an $N\times P$ data matrix with $i^{th}$ column $z_{i}$, ${\bf{y}}$ is a $N\times 1$ non-positive response vector, $\lambda_{1}\geq 0$ is the LASSO penalty, $\lambda_{2}\geq 0$ is the ridge penalty, and $\lambda_{1}+\lambda_{2}>0$. We denote the learning rate by $\alpha$, and the $i^{th}$ entry in the parameter vector after $k$ iterations as $\beta_{k}[i]$, with $10\leq i\leq P-$ 1.
@@
Our analysis relies heavily on the observation that for convex non-differentiable loss functions, the NDGM recursion in (2) effectively implements the subgradient method (Bertsekas, 1997; Shor, 2012). The theoretical properties of this method are well-studied in optimization literature, with key findings summarized in the lecture notes of Boyd et al. (2003) and Tibshirani (2015). The subgradient method and GD recursions have different properties, which play a crucial role in our discussions. We will reference relevant portions of these lecture notes as they become pertinent to our analysis.
@@
Our formulations in (1) and (2) use a constant learning rate. We explain this choice by recalling definitions from optimization literature (see Chapter 1 of Bertsekas (1997)). The training regimen in (1) is called a constant step-size regime, where GD is initialized randomly and iterated with a constant learning rate until the loss stops decreasing. Constant step-size GDs approach the stationary point but rarely converge to it unless the loss function has desirable properties like $L-smoothness$ (Proposition 1.2.3 in Bertsekas (1997)). For provable convergence in more general settings, studies often use the diminishing step-size regime, characterized by:
@@
where $\lim_{t\rightarrow\infty}\alpha_{t}=0$, and $\sum_{t}\alpha_{t}\rightarrow\infty$. Convergence results in this regime provide theoretical guarantees under ideal conditions. In practice, diminishing step-size GDs use stopping criteria to end training. We term the combination of diminishing step-size and stopping criteria commonly used in neural network literature the "reducing" step-size regime.
In this regime, training proceeds as follows: starting with weight $\beta_{0}$, GD runs for $N_{0}$ iterations with learning rate $\alpha_{0}$, followed by $N_{1}$ iterations with reduced rate $\alpha_{1}<\alpha_{0}$, and so on, until satisfactory results are achieved. This involves chaining together $T$ constant step-size GD runs, with the $(i+1)^{th}$ run initialized at $\beta_{u}$ with learning rate $\alpha_{i}$, where $u=\sum_{j=0}^{i-1}N_{j}$. While theoretical guarantees of the diminishing step-size regime may not be realized, the limiting dynamics in this regime are those of a constant step-size GD, with learning rates and weight initializations chosen from the $T^{th}$ run. Therefore, we exclusively study the constant step-size problem, assuming appropriate weight initializations and learning rates.
@@ 3 NDGM and convergence analysis
Recent theoretical work by Davis et al. (2020) has shown that in the diminishing step-size regime, NDGMs converge to a local optimal solution in the theoretical limit under fairly general conditions. However, these conditions are rarely satisfied in practice. For instance, Assumption A.3 in Davis et al. (2020) requires that the sum of the learning rates should diverge in the diminishing step size regime. While this is a standard assumption for GD analysis, it’s often violated in practice. The prevalent step-size reduction in most training scripts involves decaying the learning rate by a constant multiplicative factor at predefined intervals (e.g., StepLR in PyTorch [^2] https://rb.gy/pub4s), resulting in a convergent geometric series of learning rates, directly contradicting this assumption.
For differentiable functions, we know that even with a convergent geometric series of learning rates, the GD sequence approaches the vicinity of the optimal solution. However, it’s unclear if this holds true for NDGM sequences. More precisely, can we assert that the dynamics of NDGM and GD sequences are identical?
@@
We demonstrate the dynamics of NDGM and GD sequences are different by arriving at a contradiction. Suppose the sequence $\{\beta_{k}\}_{k=0}^{\infty}$ obtained by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})=L_{1}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1},0,0.01)$ has the same properties as a GD sequence. Then, the entire NDGM sequence can be characterized using the "Capture Theorem":
@@
Let $x_{k+1}=x_{k}-\alpha\nabla f(x_{k})$ be a sequence generated by running GD on a continuously differentiable function, $f(x)$ such that $f(x_{k+1})\leq f(x_{k})$ for all $k$. If $x^{*}$ is an isolated local minimum of $f(x)$ in a unit ball around $x^{*}$, and $||x_{0}-x^{*}||<1$, then $||x_{k}-x^{*}||<1$ for all $k>0$.
@@
Result follows by using $\bar{\epsilon}=1$ in Proposition 1.2.5 of Bertsekas (1997). ∎
Therefore, if $||\beta_{0}||_{\infty}<1$, and $L_{2}(\beta_{k+1};{\bf{Z}}_{1},{\bf{q}}_{1})\leq L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ for all $k\geq 0$, we expect $||\beta_{k}||_{\infty}<1$ for all $k\geq 0$. To validate this claim, we generate a $20\times 500$ matrix ${\bf{Z}}_{1}$ with entries uniformly sampled on $[-1,1]$, a $20\times 1$ vector ${\bf{q}_{1}}$ with entries uniformly sampled on $[-1,0]$, and the sequence $\{\beta_{t}\}_{t=0}^{\infty}$ by running NDGM on $L_{2}(\beta;{\bf{Z}}_{1},{\bf{q}}_{1})$ with entries in $\beta_{0}$ uniformly sampled on $[-1,1]$.
As shown in Figure 1, although the loss function decreases monotonically with iteration number, the infinity norm of the NDGM sequence eventually exceeds 1, contradicting our assumption. This non-capturing behavior is expected; as Boyd et al. (2003) note, "Unlike the ordinary gradient method, the subgradient method is not a descent method; the function value can (and often does) increase".
Refer to captionRefer to caption
Figure 1: Loss Function and Parameter Norm During Training. Left: Loss function $L_{2}(\beta_{k};{\bf{Z}}_{1},{\bf{q}}_{1})$ vs. iteration $k$. Right: Infinity norm $\|\beta_{k}\|_{\infty}$ vs. iteration $k$. The monotonically decreasing loss contrasts with the non-monotonic parameter norm, indicating differing dynamics between NDGM and GD.
@@
Given the distinct dynamics of GDs and NDGMs, we must reconsider the assumptions appropriate for convergence analysis in neural network training. Many analyses in the literature (Bottou et al., 2018; Zhang et al., 2022; Défossez et al., 2020; You et al., 2019) assume that the loss function $f(\beta)$ is continuously differentiable and $L-smooth$, i.e., there exists a constant $L$ such that:
@@
for all $x$, $y$ in the domain of $f$. However, this assumption fails for non-differentiable functions where the gradient doesn’t exist at various points. One might consider salvaging this assumption by limiting its validity to a local context around the local optimal solution.
Unfortunately, even this approach is untenable. As Laurent and Brecht (2018) state, "local minima of ReLU networks are generically nondifferentiable. They cannot be waved away as a technicality, so any study of the loss surface of such network must invoke nonsmooth analysis". This observation aligns with established ideas in convex optimization. For instance, the development and application of second-order cone programs (SOCPs) for convex non-smooth problems was motivated by similar considerations (Boyd et al., 2004) [^3] See 26:30 to 29:00 in https://rb.gy/ab0aeh. In both SOCPs and ReLU networks, local minima often occur at points of non-differentiability, necessitating non-smooth analysis.
Moreover, assuming $L-smoothness$ significantly overestimates the convergence speed of NDGMs. For an error level of $\epsilon$, GD applied to an $L-smooth$ convex loss function requires $O(1/\epsilon)$ steps for convergence. In contrast, NDGM applied to a convex non-differentiable loss function satisfying the Lipschitz condition:
@@
takes $O(1/\epsilon^{2})$ steps (see Section 7.1.3 of Tibshirani (2015) for the result and a visual illustration of how much slower $O(1/\epsilon^{2})$ is than $O(1/\epsilon)$). Thus, the extensive literature on GD convergence analysis for $L-smooth$ functions likely has limited relevance to the convergence dynamics of modern non-smooth neural networks. Generally, we expect smooth neural networks to converge faster than non-smooth ones.
@@
We conduct two sets of experiments on CNNs of varying depths to validate our intuition in a more general context (see the Appendix A.1 for architecture details). In each set, we perform two identical simulations for every network, with the only difference being the choice of activation function: one simulation uses the non-smooth ReLU activation function, while the other uses the smooth and asymptotically equivalent GELU activation function (Hendrycks and Gimpel, 2016). The first set of experiments excludes batch normalization, while the second includes it.
Figure 2 shows that without batch normalization, the gap in convergence speeds between smooth and non-smooth networks widens as network depth increases, aligning with our expectations. This trend is further substantiated by Hendrycks and Gimpel (2016), who demonstrate accelerated convergence in GELU networks compared to ReLU networks across diverse architectures and datasets (see Figures 2 through 7 in their paper). However, with batch normalization, the difference between ReLU and GELU training curves becomes indistinguishable, emphasizing its importance in the training process. We conclude that the choice of activation function is the primary factor contributing to the difference in convergence speeds between ReLU and GELU curves with increasing depth, as no other principled reason appears evident.
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Training curves comparing smooth and non-smooth CNNs of depth 11 (left), and 19 (right). Top: without batch normalization. Bottom: with batch normalization.
@@
The widely used $L-smoothness$ assumption, which ensures that the gradient of the loss function does not change too rapidly with respect to the parameter vector (Section 4.1 in Bottou et al. (2018)), does not hold for ReLU networks. To illustrate this, consider the 1-D case with $q(x)=\text{max}(0,x)$. In this case, $\lim_{h\rightarrow 0^{+}}(1/2h)(\nabla q(h)-\nabla q(-h))=\infty$, indicating that the gradient changes arbitrarily quickly at the origin. This necessitates the use of non-smooth analysis for studying these losses. The issue, as previously discussed, is that non-differentiable functions have kinks corresponding to non-differentiable points, and local minima frequently lie at these kinks. Therefore, gradient-based approaches developed for differentiable functions are insufficient since they cannot access these points of interest (see Appendix A.2 for geometric intuitions into this fact). The inability of gradient descent methods to effectively maneuver non-smooth functions and the associated pathologies form a recurring theme in this paper. $L-smoothness$ is one of the strongest assumptions after convexity, guaranteeing the existence and boundedness of the first and second derivatives—criteria unmet by most loss functions in neural network literature. To emphasize this point, we show in a later section that results derived for smooth networks may not generalize even to all Lipschitz continuous convex differentiable networks.